### v2

# Benedict Quartey, Ifrah Idrees, Jason Xinyu Liu, Stefanie Tellex

Abstract
Efficiently learning and executing long-horizon mobile manipulation (MoMa) tasks is crucial for advancing robotics in household and workplace settings. However, current MoMa models are data-inefficient, underscoring the need for improved models that require realistic-sized benchmarks to evaluate their efficiency, which do not exist. To address this, we introduce the LAMBDA ($\lambda$) benchmark (Long-horizon Actions for Mobile-manipulation Benchmarking of Directed Activities), which evaluates the data efficiency of models on language-conditioned, long-horizon, multi-room, multi-floor, pick-and-place tasks using a dataset of manageable size, more feasible for collection. The benchmark includes 571 human-collected demonstrations that provide realism and diversity in simulated and real-world settings. Unlike planner-generated data, these trajectories offer natural variability and replay-verifiability, ensuring robust learning and evaluation. We benchmark several models, including learning-based models and a neuro-symbolic modular approach combining foundation models with task and motion planning. Learning-based models show suboptimal success rates, even when leveraging pretrained weights, underscoring significant data inefficiencies. However, the neuro-symbolic approach performs significantly better while being more data efficient. Findings highlight the need for more data-efficient learning-based MoMa approaches. $\lambda$ addresses this gap by serving as a key benchmark for evaluating the data efficiency of those future models in handling household robotics tasks.
[…]
figureOverview. LAMBDA ($\lambda$) is a benchmark comprised of 571 language-conditioned long-horizon (room-to-room and floor-to-floor navigation) mobile manipulation (pick-and-place) tasks paired with human-collected expert demonstrations from both simulated and real-world environments. The illustrated examples in both the simulated and real-world environments span multiple rooms, which is a skill expected of daily household and workspace robots. We use a quadruped robot for its ability to recover from falls and to navigate complex terrain such as stairs, which is essential for floor-to-floor tasks.
I INTRODUCTION
Improving efficient learning for long-horizon mobile manipulation (MoMa) tasks is vital for pushing the boundaries of robotics in household and workplace applications. These tasks, such as fetching an item from a specific location or organizing objects across multiple rooms, often involve understanding language commands, navigating complex spaces, and performing precise pick-and-place actions. Although significant strides have been made in tabletop manipulation with various architectures and approaches for more data-efficient models [1, 2, 3, 4], existing MoMa models rely on large-scale datasets with extensive amounts of trajectories [5, 6, 7]. Collecting such large-scale datasets is resource-intensive and requires substantial time, manpower, and infrastructure.This data inefficiency arises from the inherent challenges of MoMa tasks: large observation and state spaces, diverse task requirements, and the need for generalization in home and workplace settings. Addressing these challenges demands the development of more data-efficient approaches for MoMa, supported by benchmarks that can guide and evaluate this progress effectively.
Existing benchmarks do not assess models for data efficiency, as they rely on unrealistically large sizes, for long-horizon, language-conditioned MoMa tasks, leaving a significant gap in evaluating and advancing progress in this area. For instance, most lack language instruction annotations for the tasks [8, 9], even though natural language provides an intuitive and accessible interface for specifying robot objectives, especially for household and workplace users. Also, most benchmarks do not support long-horizon tasks [5, 10] requiring navigation across rooms and floors, which are essential for realistic household scenarios. Furthermore, while some benchmarks support simulated environments, they often lack real-world data [11, 12], which are crucial for ensuring models’ effectiveness in practical deployments. Some current benchmark datasets also rely on planner-generated trajectories [13, 14], which lack the natural variability and expertise of human-collected demonstrations. Finally, many benchmarks lack quadruped data [15, 16], which are crucial for terrains like stairs for floor-to-floor navigation, and maintaining stability during tasks.
To address these challenges, we propose the LAMBDA ($\lambda$) benchmark (Long-horizon Actions for Mobile-manipulation Benchmarking of Directed Activities). $\lambda$ evaluates models on a realistic amount of demonstrations of language-conditioned, multi-room, multi-floor, pick-and-place tasks requiring long-horizon planning and execution, addressing the limitations of existing benchmarks. The dataset consists of 571 high-quality trajectories, which were crowdsourced and manually collected to ensure realism, diversity, and replay-verifiability. Crowdsourcing was specifically conducted to capture natural, flowing instructions rather than templated commands, providing a realistic distribution of tasks that people want robots to perform in household and workplace environments. By including language-conditioned tasks, $\lambda$ supports intuitive and accessible interactions, making it practical for real-world users. Unlike planner-generated data, human-collected trajectories in $\lambda$ provide natural variability and expertise, better capturing real-world robot operations. Additionally, $\lambda$ supports long-horizon tasks involving navigation across rooms and floors, reflecting the realistic demands of indoor scenarios. Its simulated environments enable diverse testing scenarios and object variability, ensuring flexibility and standardization, while real-world data ensures models are prepared for practical deployment. Finally, the inclusion of armed quadrupeds enables evaluation on complex terrains, such as stairs, highlighting their stability and adaptability in challenging conditions.
We used $\lambda$ to benchmark several models and establish baseline performance for long-horizon MoMa tasks. Behavior cloning (BC) models trained on the dataset showed low success rates, $2.7\%$ and $5.1\%$, indicating data inefficiency while confirming the dataset’s ability to support learning. A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved an $44.4\%$ success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems in the future. These results demonstrate $\lambda$’s effectiveness in highlighting data efficiency limitations of current models and guiding the future development of improved architectures for indoor robotics.
Our contributions are summarized as follows:
  • $\lambda$ benchmark: We introduce a novel benchmark specifically tailored in size to evaluate data efficiency in MoMa, emphasizing language-conditioned, long-horizon pick-and-place indoor tasks in realistic household and workplace settings.
  • High-quality dataset: $\lambda$’s tasks are paired with 571 human-collected trajectories, providing natural variability, replay-verifiability, and a balance of simulated and real-world data for robust model evaluation.
  • Comprehensive evaluation: We benchmark several models, including behavior cloning and neuro-symbolic modular approaches, providing a baseline for future model development for MoMa tasks of this nature.
Benchmark Natural Language Long-Horizon Demos Human-Collected Sim Real Quadruped
SafeAgentBench
HumanoidBench
Mobile ALOHA
BEHAVIOR-1K
ManiSkill-HAB
InfiniteWorld
RoboSuite
RoboCasa Some Some
BRMData
DeMoBot
MoMaRT
M3Bench
BiGym
RT-1
$\lambda$
TABLE I: Comparison. Different mobile manipulation benchmarks. $\lambda$ is the first mobile manipulation benchmark that comprises human-collected demonstrations of language-conditioned long-horizon tasks in simulated and real-world environments utilizing a quadruped. “Long-Horizon,” in this case, is defined as long-horizon indoor navigation across multiple rooms and/or floors. “Some” indicates that only a portion of the tasks/demonstrations possess the respective quality or attribute.
II Related Work
As robot learning algorithms evolve rapidly, benchmarks have become critical for evaluating algorithmic decisions and ensuring consistent performance comparisons in standardized scenarios. However, existing benchmarks, shown in Table I, fail to adequately evaluate models for data efficiency in long-horizon, language-conditioned indoor MoMa tasks, leaving significant gaps in advancing this area. For example, most benchmarks [8, 9, 17, 14, 18, 19, 20, 16, 10], lack natural language annotations, despite its importance as an intuitive interface for specifying tasks, particularly in household and workplace scenarios. Language conditioning allows robots to interpret human commands directly, making this an essential feature for realistic applications.
Moreover, almost all benchmarks [8, 9, 14, 18, 15, 19, 20, 16, 13, 10, 5], do not support long-horizon tasks involving navigation across rooms and floors. These tasks are critical for indoor robotics, where robots must traverse long distances to perform daily operations. Without this capability, benchmarks fail to reflect the complexity of real-world scenarios. While most benchmarks provide simulated environments, Mobile ALOHA [9] and BRMData [19] lack simulation data entirely. Simulated environments offer diverse scenes and objects for flexible testing across varied scenarios and provide a standardized, portable testbed that can be easily used by others due to their fixed conditions. On the other hand, most lack real-world data [11, 8, 14, 12, 18, 16, 13, 10], which is essential for providing a benchmark grounded in realistic scenarios where robots will ultimately be deployed, ensuring models are tested under practical and uncontrolled conditions.
Most benchmarks provide expert demonstrations for the tasks, which are crucial for establishing ground-truth data to train learning-based methods and test models. However, a few, SafeAgentBench [11], HumanoidBench [8], BEHAVIOR-1K [17], and InfiniteWorld [12], lack this attribute, limiting their utility for model training and evaluation. M3Bench [13] and ManiSkill-HAB [14] depend on planner-generated trajectories. These lack the natural variability and expertise of human-collected demonstrations, which more accurately reflect real-world conditions and improve the quality of learned behaviors by providing nuanced, high-quality examples that enable better generalization.
Finally, most benchmarks from Table I, with the exceptions of RoboCasa [15], RoboSuite [18], and DeMoBot [20], do not feature armed quadruped data, which is essential for tasks requiring stability and adaptability. Quadrupeds with manipulators can handle uneven terrains, such as stairs, while carrying objects, making them ideal for multi-floor environments, such as the ones present in $\lambda$. Their robustness as quadrupeds allows them to maintain stability in obstacle-dense environments, withstand collisions, and recover by picking themselves up if they fall, making them especially suitable for realistic and demanding scenarios. The absence of this embodiment limits the applicability of the vast majority of benchmarks to realistic robotic use cases.
By addressing these limitations, the $\lambda$ benchmark provides a comprehensive benchmark to test the data efficiency of MoMa models. It integrates language-conditioned tasks, long-horizon trajectories spanning multiple rooms and floors, and both simulated and real-world data. Additionally, its human-collected demonstrations ensure realistic variability, and its inclusion of armed quadruped data makes it uniquely suited for advancing robotics in complex real-world settings.
III $\lambda$ Benchmark
We present LAMBDA ($\lambda$), a benchmark designed to evaluate model data efficiency on long-horizon MoMa tasks. $\lambda$ features language-conditioned, multi-room, multi-floor pick-and-place tasks that span a wide range of diverse environments across both simulated and real settings. The tasks include their respective human-collected demonstrations, capturing natural variability and providing realistic scenarios for training and evaluation. This section details the unique challenges of the $\lambda$ tasks (see Section III-A), the simulation dataset (see Section III-B), and the real-world dataset (see Section III-C).
III-A Challenges of $\lambda$ Tasks
Data Efficiency in MoMa. $\lambda$ is designed to test the data efficiency of learning-based models by offering a dataset of moderate size, deliberately curated to strike a balance between realism and practicality. While recent data-efficient tabletop manipulation models [1, 2, 3, 4] often rely on double-digit demonstrations, we provide a few hundred demonstrations, recognizing the increased complexity of MoMa. MoMa encompasses not only higher-dimensional observation and action spaces but also significantly larger and more varied task spaces. This deliberate dataset size ensures sufficient diversity for learning while maintaining a manageable scale that avoids the excessive computational and collection costs of unrealistically large datasets.
Complex Spaces. The inclusion of MoMa tasks introduces significant complexity due to the interplay of action, observation, state, and task spaces. Unlike fixed single-arm setups, MoMa requires navigating and interacting within dynamic three-dimensional environments, often spanning multiple rooms or floors. This necessitates the seamless integration of high-level planning and low-level motion control, creating challenges for robotic models to operate effectively. For example, consider a household task where the agent must navigate to a dining table, pick up a plate, and transport it to a countertop in another room. Even in this straightforward pick-and-place task, the complexity of the action, state, and observation spaces introduces substantial variability. The agent might approach the table from different directions, choose from various grasping points on the plate, or follow alternative routes to the destination. This multi-modal nature means that the same task can be solved in multiple ways, further increasing the demands on models to generalize across diverse strategies and environmental configurations. Moreover, the complexity of these integrated spaces enables a broader range of tasks, from transporting objects across floors to rearranging items in dynamic environments. For our benchmark, we focus on pick-and-place tasks, as they represent a significant portion of daily activities and provide a foundational challenge for MoMa models to address.
Long-Horizon Tasks. A key challenge in $\lambda$ lies in its dedicated emphasis on long-horizon tasks, where navigation spans multiple rooms and floors in indoor environments such as households and workplaces. These tasks inherently demand long-term planning, often requiring hierarchical strategies that integrate high-level task objectives with low-level motion execution. The agent must account for spatial relationships, sequential task dependencies, and the compounding effects of errors over time. Error propagation is a critical challenge across many learning paradigms, exemplified by covariate shift in imitation learning. For instance, navigating to a distant room to retrieve an object and returning it to another location involves traversing obstacles and staircases while adhering to task objectives and maintaining situational awareness. Such tasks are representative of real-world scenarios where robots are expected to perform complex, multi-step instructions over extended periods.
Realistic Task Distribution. To ensure the tasks in $\lambda$ reflect practical needs, we used crowdsourcing to gather a diverse and realistic set of tasks that people expect robots to perform in household and workplace settings. This method avoided templated task specifications by capturing natural, human-generated instructions. The resulting dataset encompasses a variety of pick-and-place MoMa tasks, aligning with real-world expectations and exposing models to inherently varied and context-dependent scenarios.
III-B Simulated Dataset
The dataset for our simulated tasks consists of 521 trajectories collected across 20 rooms in five distinct environments, designed to ensure diversity in both object types and room layouts. We utilized the AI2THOR [21] simulator and specifically employed RoboTHOR [22] environments, which feature multiple interconnected rooms suitable for long-horizon tasks.
For robot control, we used ManipulaTHOR [23], which includes an arm capable of fine manipulation and provides low-level pose data. Each trajectory includes a range of attributes, including but not limited to RGB-D egocentric observations, segmentations, the poses of the robot body, end-effector, and grasped objects, as well as discrete actions taken during the tasks.
To generate realistic and diverse instructions, we leveraged Prolific111prolific.com for crowdsourcing to gather 521 natural language commands from 41 participants, reflecting tasks that people would naturally expect robots to perform in indoor settings. Using these commands, we teleoperated the robot in the simulator with a custom keyboard-based teleoperation system we developed, ensuring accuracy and consistency throughout the data collection process. After teleoperation, each trajectory was replay-verified to confirm correctness and eliminate errors, ensuring high-quality data integrity.
Figure 1: Trajectories. Two mobile manipulation tasks specified via natural language along with their expert trajectories. The top orange row shows the teleoperation of the Spot robot, and the bottom orange row depicts the Spot’s egocentric observation.
III-C Real-World Dataset
The dataset for our real-world tasks comprises 50 trajectories collected across 10 rooms in three distinct environments. These environments were chosen for their large size, structural features that allow for the creation of long-horizon tasks, and their status as workplaces where robots can be particularly useful for daily activities. The first environment is a three-room laboratory, the second is a floor in a university building, and the third spans two adjacent floors connected by stairs within the same building. These spaces include kitchens, furnished lobbies, a classroom, and a staircase, offering diverse objects and room layouts to enhance the generalizability of trained models.
To handle complex terrain and ensure stability, we selected the Spot [24] quadruped mobile manipulator, equipped with an arm for pick-and-place tasks. The quadruped’s design enables it to navigate stairs and maintain balance during task execution, making it ideal for the challenging settings of our chosen environments.
We teleoperated the Spot robot using a tablet equipped with thumbsticks to collect task trajectories based on 50 natural language commands provided by seven participants. The commands, distributed as 20, 15, and 15 across the three environments, specified long-horizon room-to-room and floor-to-floor pick-and-place tasks that reflect practical scenarios in workplace settings. Task trajectories were collected at 3 Hz to ensure detailed and realistic data.
IV Evaluation
IV-A Baseline Models
To evaluate the $\lambda$ and establish baselines for future benchmarking, we test both learning-based and neuro-symbolic models. Given the growing popularity of imitation learning (IL), particularly BC, we benchmark a state-of-the-art (SOTA) BC model. These models provide insight into how data-efficient current BC algorithms are by being tested on $\lambda$, validating that the dataset supports meaningful learning by demonstrating that models can achieve measurable yet limited success. Additionally, we benchmark a neuro-symbolic system that integrates foundation models with task and motion planning (TAMP). This system, evaluated zero-shot without requiring training, investigates whether non-IL methods can achieve greater data efficiency while achieving comparable or superior performance, offering guidance for future development of data-efficient MoMa methodologies.
RT-1. Robotics Transformer (RT-1) [5] is a BC model designed for MoMa, leveraging a Transformer architecture to process natural language instructions and RGB image inputs into discretized actions. Its training on MoMa tasks, along with its status as an open-source, SOTA learning-based MoMa model, makes it particularly suitable for benchmarking on $\lambda$ as a baseline.
To adapt RT-1 for the ManipulaTHOR robot, we modified its action space to align with the robot’s capabilities. For task navigation, the robot base operates using discrete actions, while continuous deltas are utilized for body rotation and manipulation to achieve precise control. Predicting deltas between timesteps rather than absolute states improved learning stability and control smoothness, while the discretization of continuous action outputs into 256 uniformly distributed bins enabled effective representation and improved learning.
LIMP. Language Instruction grounding for Motion Planning (LIMP) [25] is a neuro-symbolic system for pick-and-place MoMa that integrates Linear Temporal Logic (LTL), foundation models, and novel TAMP algorithms to process natural language instructions into executable TAMP plans. By translating natural language instructions into LTL formulas enriched with Composable Referent Descriptors, LIMP dynamically generates context-aware specification translations and compiles them into finite-state automata for verifiably correct execution. Unlike learning-based methods, LIMP requires no training, allowing it to be evaluated zero-shot. This enables a comparison of its data efficiency against BC models, providing insight into whether such systems are a promising direction for developing more data-efficient approaches to mobile manipulation.
Scene 1 Scene 2 Scene 3 Scene 4 Scene 5 Task Gen
Train Val Test Train Val Test Train Val Test Train Val Test Train Val Test Train Val Test
Demos 330 83 108 340 86 95 337 85 99 324 82 115 333 84 104 416 51 54
Scenes 4 4 1 4 4 1 4 4 1 4 4 1 4 4 1 5 5 5
TABLE II: Data Splits. Splitting of the simulation dataset of 521 demonstrations into training, validation, and held-out test subsets for the Scene Generalization and Task Generalization experiments. For Scene Gen., a stratified sampling approach was employed to account for variations in the number of demonstrations per scene (details in Section IV-B). The “Scenes” row indicates the number of scenes contributing to each subset of demonstrations.
IV-B Experiments.
Data Split & Generalization. To evaluate generalization capabilities, we conduct experiments using carefully designed data splits for scene generalization and task generalization on the simulation dataset, highlighted in Table II.
For scene generalization, we employ a 5-fold cross-validation approach, where each of the five scenes in the dataset is held out as the test scene in turn. The held-out test set consists of all trajectories from the designated scene, while the training and validation sets are derived from the remaining four scenes. Specifically, 80% of the trajectories from each of the four remaining scenes are included in the training set, and 20% are allocated to the validation set, following a stratified split to preserve data distribution. During each fold, we train and hyperparameter tune a separate model specific to the held-out test scene. This process ensures that we assess the model’s ability to generalize to entirely unseen environments while maintaining robust experimental controls. The rotation of held-out scenes allows for a comprehensive evaluation of scene generalization across the entire dataset.
For task generalization, the data split is consistent across all five scenes. The training set comprises 80% of the trajectories from each scene, the validation set includes 10%, and the test set includes the remaining 10%. Unlike scene generalization, the held-out test set in this experiment evaluates the model’s ability to generalize to novel tasks rather than unseen environments. Hyperparameter tuning is performed, resulting in a single model optimized for task generalization on the test set.
Training. To benchmark the data efficiency of modern BC methods, we train the RT-1 model from scratch without any pretrained parameters on the simulation tasks. This approach evaluates how well these models and their architectures learn from our modestly sized yet realistic $\lambda$ dataset, providing a measure of their data efficiency specifically for diverse and long-horizon MoMa tasks. Both models are trained using the scene generalization and task generalization experiments, ensuring that their capabilities are evaluated under varying conditions.
Fine-tuning. To further evaluate data efficiency, we fine-tune RT-1 by loading pretrained parameters and testing its performance on the $\lambda$ benchmark simulation tasks. We investigate whether pretrained knowledge helps or hurts model generalization on our tasks. Empirical analysis showed that freezing some model layers led to better performance than not freezing any layers, so this approach was used for the final experiment. Using this setup, we conducted the task generalization experiment to measure how much pretrained MoMa knowledge impacts RT-1’s ability to adapt efficiently to new MoMa tasks.
V Discussion
V-A Results
Scene Generalization. The success rates for the Scene Generalization experiment, detailed in Table III, are exceptionally low, with an overall average of just $2.7\%$ across all scenes. None of the individual scenes achieved success rates exceeding $5\%$, indicating significant challenges in generalizing to novel scenes with unseen room layouts, furniture configurations, and object placements for long-horizon MoMa tasks.
Scenes 1 and 3 exhibited relatively higher success rates compared to the others. This can be attributed to a relatively high proportion of evaluation tasks in these scenes involving objects that were positioned directly in front of or adjacent to the robot’s initial spawn location, reducing the complexity of navigation and manipulation required. In contrast, Scene 1 exhibited the lowest performance, likely due to the robot spawning in an enclosed and isolated room where many tasks necessitated navigating out of the room to retrieve the target object, significantly increasing task difficulty.
Task Generalization. The success rates for the Task Generalization experiment for RT-1, as presented in Table IV, are low overall, though they are slightly higher compared to the average performance in the Scene Generalization experiment. This improvement is likely attributed to the model having access to all the scenes during training, enabling it to better generalize to new tasks within already familiar environments.
We also evaluated on seen tasks, which were uniformly sampled from the training data to match the size of the unseen data from the test split to measure performance on previously encountered tasks. Surprisingly, RT-1 achieved a marginally higher success rate on unseen tasks ($5.1\%$) compared to seen tasks ($4.6\%$). This indicates that the model’s training did not overfit to the training data, which is a positive outcome. One plausible explanation is that the sampler may have inadvertently selected a disproportionate number of especially challenging tasks for the seen evaluation set, rendering them overly difficult for the model despite having been part of the training data.
Task Generalization with Fine-Tuning. To evaluate whether pretrained parameters improve generalization on our benchmark tasks, we fine-tuned RT-1 with some of its layers frozen and tested it on both seen and unseen tasks. Interestingly, the unseen task performance was marginally lower compared to the fully trained model’s performance on unseen tasks, as seen in Table IV. This result suggests that the pretrained weights may have skewed the model’s representation toward the original training tasks and away from effectively adapting to the unique challenges of our benchmark, especially since the original was trained in the real world
LIMP. The results presented thus far indicate that learning-based models, such as RT-1, face significant challenges in addressing the complexities of long-horizon MoMa tasks. To explore an alternative paradigm, we benchmarked the LIMP system on the Task Generalization experiment.
The evaluation was conducted in real-world environments using Spot, assessing 45 tasks, including a mix of tasks originally specified for both simulated and real-world environments. Since LIMP was originally designed for real-world execution, this mixed evaluation approach was necessary to also test its performance on simulation-specified tasks. For these simulation-designed tasks, real-world setups were constructed to approximate their simulated counterparts. This setup enabled us to evaluate LIMP’s versatility across both task types, providing insights into its adaptability to different environmental contexts. As shown in Table IV, LIMP achieved a relatively impressive average success rate of $44.4\%$ on all novel tasks, vastly outperforming RT-1 without requiring any training, making it relatively more data efficient.
One key factor contributing to LIMP’s success is its use of pre-generated maps of the environments. These maps provide the robot with precise information about the absolute and relative locations of objects and regions, eliminating the need to search and enabling it to interpret the fine-grained semantics specified in language-conditioned tasks. Additionally, the maps allow the robot to utilize robust motion planning algorithms for navigation, which are less susceptible to covariate shift compared to the mapless autoregressive action decoding used in BC models like RT-1. This robustness is particularly critical for the long-horizon navigation aspect of the tasks, where the inherent large observation and state spaces exacerbate the challenges faced by learning-based methods.
Another significant advantage of LIMP is its incorporation of multimodal foundation models, which endow it with extensive prior knowledge to reason about task requirements and perceptual inputs in a zero-shot manner. This capability enables LIMP to handle the semantic and contextual complexities of our benchmark tasks without requiring task-specific training. While LIMP demonstrates considerable promise, it still falls short of perfect task execution, indicating room for improvement in handling the intricacies of our long-horizon MoMa tasks. Overall, LIMP represents a promising direction for future system development, blending the strengths of learning-based approaches with those of classical planning.
V-B Ablations
Impact of Dataset Size. The poor performance observed with learning-based methods is expected to stem from the limited, realistically sized dataset of the $\lambda$ benchmark, but this needs to be further examined. Will simply scaling the data improve the performance? Will these models effectively and efficiently use more data to learn better? To address this, we investigate how performance scales when progressively larger subsets of the dataset ($25\%$, $50\%$, $75\%$, and $100\%$) are used for training. This evaluates whether these models can effectively utilize limited data or require substantially more, directly testing their data efficiency and scalability for long-horizon MoMa tasks.
The results from the Task Generalization setup, summarized in Table V, surprisingly reveal no clear upward trend in performance gains despite progressively increasing the training data for RT-1 and hyperparameter tuning per subset. We attribute part of the variation to the inherent complexity of these long-horizon MoMa tasks and the variability in the sampled subsets, where smaller subsets can incidentally include a higher proportion of comparatively easier scenarios. Overall, this suggests that simply providing more samples does not guarantee better learning outcomes for these tasks, and it is likely that limited data is not the primary cause of the poor performance. Moreover, since the linear increase in data size did not result in proportional improvement, it indicates that the model is unable to efficiently learn from the additional data provided, deeming it data inefficient. This raises questions about deeper limitations, such as architectural constraints, warranting further investigation.
Scene Gen
Model Scene 1 Scene 2 Scene 3 Scene 4 Scene 5 Avg
RT-1 $4.9\pm 9.9$ $0.78\pm 4.4$ $4.3\pm 9.4$ $2.0\pm 6.7$ $1.4\pm 5.8$ $2.7$
TABLE III: Scene Generalization results. Success rates for the Scene Generalization experiment on models trained from scratch. The success rate is computed by averaging the scores of four subtasks per task, with each successful subtask contributing one point, and then averaging across all tasks. The subtasks are: navigation to the target object, grasping the target object, navigation to the goal with the target object, and placing the target object at the goal. $\pm$ indicates the standard deviation across all evaluated tasks. “Avg” indicates the average of all the scenes.
Task Gen Task Gen FT
Model Seen Unseen Seen Unseen
RT-1 $4.6\pm 9.7$ $5.1\pm 10.1$ $4.6\pm 9.7$ $4.6\pm 9.7$
LIMP $44.4\pm 49.7^{*}$
TABLE IV: Task Generalization results. Success rates for Task Generalization experiments, encompassing both training from scratch and fine-tuning of pretrained parameters. The computation is the same as Table III. “Seen” refers to evaluation on the training subset, while “Unseen” refers to evaluation on the held-out test subset. Denotes zero-shot evaluation with no training on a mix of simulation and real-world tasks.
Impact of Architecture. To determine if data inefficiency stems from the architecture used, we conduct a model comparison experiment by evaluating a different architecture. We developed a custom LSTM-based model by replacing RT-1’s Transformer with an LSTM while retaining the same encoders for a consistent and isolated comparison, as LSTMs have been previously used for sequential robotic tasks. The LSTM variant is evaluated under the task generalization setup using subset experiments with $25\%$, $50\%$, $75\%$, and $100\%$ of the dataset, enabling us to analyze how its performance scales with increasing data. By comparing the full dataset results and the performance rate of change (RoC) against RT-1, we aim to determine whether data inefficiency is tied to architectural limitations or if the issue lies elsewhere.
As shown in Table V, the LSTM variant exhibits no upward trend when moving from $25\%$ to $100\%$ of the data, indicating it is also data inefficient. To compare the trends between the two architectures, we analyze their average rates of change. Both exhibit similarly non-positive trends, indicating that the data inefficiency is unlikely to primarily stem from the choice of architecture and seems to span across different architectures, since they are not efficiently making use of more data. Furthermore, since RT-1’s $100\%$ performance is somewhat better than the LSTM variant’s, and its RoC is slightly higher, it indicates marginally better data efficiency, further supporting the argument that architecture is unlikely the primary cause of RT-1’s data inefficiency observed in the main experiments.
Since architecture does not appear to be the primary issue, it is plausible that these long-horizon indoor MoMa tasks are inherently complex and challenging for learning-based methods. Identifying the exact causes of data inefficiency in these models requires further investigation, which is beyond the scope of this paper and is left for the community to explore in future work.
VI Conclusion
We introduce LAMBDA ($\lambda$), a challenging benchmark consisting of a realistic amount, 571, of language-conditioned, long-horizon mobile manipulation tasks, spanning room-to-room and floor-to-floor navigation in indoor environments. Each task is paired with human-collected expert demonstrations from both simulated and real-world settings.
To evaluate the benchmark, we tested multiple models, including RT-1 and LIMP, to assess their performance and data efficiency on these complex tasks. Learning-based models such as RT-1 demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to $\lambda$’s long-horizon tasks. In contrast, the neuro-symbolic system LIMP significantly outperformed RT-1, showcasing its potential as a promising direction for the development of future mobile manipulation systems capable of addressing the challenges posed by $\lambda$’s tasks. We hope that this benchmark will serve as a valuable resource for the robotics community, fostering the development of more advanced models and systems that can efficiently address daily household and workplace tasks.
Limitations and Future Work. While the inclusion of a quadruped like Spot in the real world tasks highlights the advantages of quadrupedal platforms, the lack of simulated quadruped data limits the ability to leverage diverse environments that simulators can provide, such as a broader range of floor-to-floor trajectories. Expanding the benchmark to include simulated quadruped tasks would enable greater generalization and provide richer datasets for evaluation. Furthermore, extending the benchmark to include tasks beyond pick-and-place, such as those involving more intricate manipulation, would increase its relevance and applicability to a wider range of robotic systems and models. Future work could also benchmark reinforcement learning models, which, while not the focus of this study, are known for their sample inefficiency. Evaluating their performance on $\lambda$’s sparse rewards would offer additional insights into their suitability for long-horizon mobile manipulation tasks and broaden the scope of the benchmark.
ACKNOWLEDGMENT
This work is supported by ONR under grant award numbers N00014-22-1-2592 and N00014-23-1-2794, NSF under grant award number CNS-2150184, and with support from Amazon Robotics. We also thank Aryan Singh, George Chemmala, Ziyi Yang, David Paulius, Ivy He, Lakshita Dodeja, Mingxi Jia, Benned Hedegaard, Thao Nguyen, Selena Williams, Tuluhan Akbulut, Chaerin Min, and George Konidaris for their help in various phases of work.
Model $25\%$ $50\%$ $75\%$ $100\%$ Avg RoC
RT-1 $8.3\pm 11.8$ $8.9\pm 12.0$ $4.8\pm 11.2$ $5.1\pm 10.1$ $-0.04$
LSTM $8.3\pm 11.8$ $2.7\pm 7.7$ $5.4\pm 10.3$ $1.4\pm 5.7$ $-0.092$
TABLE V: Ablations results. Success rates for Task Generalization on growing subsets of the simulation dataset for both RT-1 and its LSTM variant. The computation is the same as Table III. “Avg RoC” indicates the average rate of change between subset results.

### v3

@@
 We used $\lambda$ to benchmark several models and establish baseline performance for long-horizon MoMa tasks.
-Behavior cloning (BC) models trained on the dataset showed low success rates, $2.7\%$ and $5.1\%$, indicating data inefficiency while confirming the dataset’s ability to support learning.
+Behavior cloning (BC) models trained on the dataset showed low success rates, such as $2.7\%$ and $5.1\%$, indicating data inefficiency while confirming the dataset’s ability to support learning.
 A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved an $44.4\%$ success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems in the future.
@@
 To evaluate the $\lambda$ and establish baselines for future benchmarking, we test both learning-based and neuro-symbolic models.
-Given the growing popularity of imitation learning (IL), particularly BC, we benchmark a state-of-the-art (SOTA) BC model.
+Given the growing popularity of imitation learning (IL), particularly BC, we benchmark two state-of-the-art (SOTA) BC models.
 These models provide insight into how data-efficient current BC algorithms are by being tested on $\lambda$, validating that the dataset supports meaningful learning by demonstrating that models can achieve measurable yet limited success.
@@
 Predicting deltas between timesteps rather than absolute states improved learning stability and control smoothness, while the discretization of continuous action outputs into 256 uniformly distributed bins enabled effective representation and improved learning.
+MotionGlot.
+MotionGlot (MG) [25] is designed to address the problem of varying action and observation dimensions in multi-embodiment models [5, 26].
+Leveraging training principles from language models (LLMs), MG proposes a training procedure for motion generation across multiple embodiments as a next-token prediction problem.
+Here we adapt MG for the MoMa tasks, given a set of demonstration trajectories with image observation ($o_{i}$) action ($a_{i}$) pairs in the form of $\tau=\{(o_{0},a_{0}),(o_{1},a_{1}),...,(o_{T},a_{T}))\}$, MG learns a model to map a given observation to action.
+This process involves two stages: First, a VQ-VAE is employed to learn a discrete latent codebook for the images, forming the observation vocabulary, while discrete binning is applied to tokenize the actions.
+In the second stage, both the observation and action vocabularies are integrated with the text vocabulary from GPT-2 [27].
+This unified vocabulary allows for action generation similar to text generation.
+Finally, the instruction-tuning template from MG is used for text command-based action generation.
 LIMP.
@@
 Splitting of the simulation dataset of 521 demonstrations into training, validation, and held-out test subsets for the Scene Generalization and Task Generalization experiments.
-For Scene Gen., a stratified sampling approach was employed to account for variations in the number of demonstrations per scene (details in Section IV-B).
 The “Scenes” row indicates the number of scenes contributing to each subset of demonstrations.
@@
 Task Generalization.
-The success rates for the Task Generalization experiment for RT-1, as presented in Table IV, are low overall, though they are slightly higher compared to the average performance in the Scene Generalization experiment.
+The success rates for the Task Generalization experiment for RT-1 and MotionGlot, as presented in Table IV, are low overall, though they are slightly higher compared to the average performance in the Scene Generalization experiment.
 This improvement is likely attributed to the model having access to all the scenes during training, enabling it to better generalize to new tasks within already familiar environments.
+Additionally, while it is unsurprising that RT-1 outperformed MotionGlot, it is notable that the margin was relatively small, particularly given that MotionGlot was originally designed for motion generation rather than specifically for mobile manipulation.
 We also evaluated on seen tasks, which were uniformly sampled from the training data to match the size of the unseen data from the test split to measure performance on previously encountered tasks.
@@
 LIMP.
-The results presented thus far indicate that learning-based models, such as RT-1, face significant challenges in addressing the complexities of long-horizon MoMa tasks.
+The results presented thus far indicate that learning-based models, such as RT-1 and MotionGlot, face significant challenges in addressing the complexities of long-horizon MoMa tasks.
 To explore an alternative paradigm, we benchmarked the LIMP system on the Task Generalization experiment.
@@
 The subtasks are: navigation to the target object, grasping the target object, navigation to the goal with the target object, and placing the target object at the goal.
-$\pm$ indicates the standard deviation across all evaluated tasks.
-“Avg” indicates the average of all the scenes.
 Task Gen Task Gen FT
@@
 RT-1 $4.6\pm 9.7$ $5.1\pm 10.1$ $4.6\pm 9.7$ $4.6\pm 9.7$
+MotionGlot $3.24\pm 8.4$ $2.4\pm 7.2$
 LIMP $44.4\pm 49.7^{*}$
@@
 Success rates for Task Generalization experiments, encompassing both training from scratch and fine-tuning of pretrained parameters.
-The computation is the same as Table III.
 “Seen” refers to evaluation on the training subset, while “Unseen” refers to evaluation on the held-out test subset.
@@
 To determine if data inefficiency stems from the architecture used, we conduct a model comparison experiment by evaluating a different architecture.
-We developed a custom LSTM-based model by replacing RT-1’s Transformer with an LSTM while retaining the same encoders for a consistent and isolated comparison, as LSTMs have been previously used for sequential robotic tasks.
+While MotionGlot was already benchmarked and provides some architectural comparison with RT-1, it is not sufficiently different since it also relies on a transformer architecture.
+To explore a more distinct alternative, we developed a custom LSTM-based model by replacing RT-1’s Transformer with an LSTM while retaining the same encoders for a consistent and isolated comparison, as LSTMs have been previously used for sequential robotic tasks.
 The LSTM variant is evaluated under the task generalization setup using subset experiments with $25\%$, $50\%$, $75\%$, and $100\%$ of the dataset, enabling us to analyze how its performance scales with increasing data.
@@
 Each task is paired with human-collected expert demonstrations from both simulated and real-world settings.
-To evaluate the benchmark, we tested multiple models, including RT-1 and LIMP, to assess their performance and data efficiency on these complex tasks.
-Learning-based models such as RT-1 demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to $\lambda$’s long-horizon tasks.
-In contrast, the neuro-symbolic system LIMP significantly outperformed RT-1, showcasing its potential as a promising direction for the development of future mobile manipulation systems capable of addressing the challenges posed by $\lambda$’s tasks.
+To evaluate the benchmark, we tested multiple models: RT-1, MotionGlot, and LIMP, to assess their performance and data efficiency on these complex tasks.
+Learning-based models (RT-1 and MotionGlot) demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to $\lambda$’s long-horizon tasks.
+In contrast, the neuro-symbolic system LIMP significantly outperformed RT-1 and MotionGlot, showcasing its potential as a promising direction for the development of future mobile manipulation systems capable of addressing the challenges posed by $\lambda$’s tasks.
 We hope that this benchmark will serve as a valuable resource for the robotics community, fostering the development of more advanced models and systems that can efficiently address daily household and workplace tasks.
@@
 Evaluating their performance on $\lambda$’s sparse rewards would offer additional insights into their suitability for long-horizon mobile manipulation tasks and broaden the scope of the benchmark.
-ACKNOWLEDGMENT
-This work is supported by ONR under grant award numbers N00014-22-1-2592 and N00014-23-1-2794, NSF under grant award number CNS-2150184, and with support from Amazon Robotics.
-We also thank Aryan Singh, George Chemmala, Ziyi Yang, David Paulius, Ivy He, Lakshita Dodeja, Mingxi Jia, Benned Hedegaard, Thao Nguyen, Selena Williams, Tuluhan Akbulut, Chaerin Min, and George Konidaris for their help in various phases of work.
 Model $25\%$ $50\%$ $75\%$ $100\%$ Avg RoC

### v5

@@
-Benedict Quartey, Ifrah Idrees, Jason Xinyu Liu, Stefanie Tellex
+Anneke Wernerfelt, Benedict Quartey, Ifrah Idrees, Jason Xinyu Liu, Stefanie Tellex
 Abstract
@@
 These tasks, such as fetching an item from a specific location or organizing objects across multiple rooms, often involve understanding language commands, navigating complex spaces, and performing precise pick-and-place actions.
-Although significant strides have been made in tabletop manipulation with various architectures and approaches for more data-efficient models [1, 2, 3, 4], existing MoMa models rely on large-scale datasets with extensive amounts of trajectories [5, 6, 7].
+Although significant strides have been made in tabletop manipulation with various architectures and approaches for more data-efficient models [1, 2, 3], existing MoMa models rely on large-scale datasets with extensive amounts of trajectories [4, 5, 6].
 Collecting such large-scale datasets is resource-intensive and requires substantial time, manpower, and infrastructure.This data inefficiency arises from the inherent challenges of MoMa tasks: large observation and state spaces, diverse task requirements, and the need for generalization in home and workplace settings.
@@
 A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved an $44.4\%$ success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems in the future.
-These results demonstrate $\lambda$’s effectiveness in highlighting data efficiency limitations of current models and guiding the future development of improved architectures for indoor robotics.
+These results demonstrate $\lambda$’s effectiveness in highlighting data efficiency limitations of current models and guiding the future development of improved ones for indoor robotics.
 Our contributions are summarized as follows:
@@
   • High-quality dataset: $\lambda$’s tasks are paired with 571 human-collected trajectories, providing natural variability, replay-verifiability, and a balance of simulated and real-world data for robust model evaluation.
-  • Comprehensive evaluation: We benchmark several models, including behavior cloning and neuro-symbolic modular approaches, providing a baseline for future model development for MoMa tasks of this nature.
+  • Comprehensive evaluation: We benchmark several models, including BC and neuro-symbolic modular approaches, providing a baseline for future model development for MoMa tasks of this nature.
 Benchmark Natural Language Long-Horizon Demos Human-Collected Sim Real Quadruped
@@
 The tasks include their respective human-collected demonstrations, capturing natural variability and providing realistic scenarios for training and evaluation.
-This section details the unique challenges of the $\lambda$ tasks (see Section III-A), the simulation dataset (see Section III-B), and the real-world dataset (see Section III-C).
 III-A Challenges of $\lambda$ Tasks
@@
 $\lambda$ is designed to test the data efficiency of learning-based models by offering a dataset of moderate size, deliberately curated to strike a balance between realism and practicality.
-While recent data-efficient tabletop manipulation models [1, 2, 3, 4] often rely on double-digit demonstrations, we provide a few hundred demonstrations, recognizing the increased complexity of MoMa.
-MoMa encompasses not only higher-dimensional observation and action spaces but also significantly larger and more varied task spaces.
+While recent data-efficient tabletop manipulation models [1, 2, 3] often rely on double-digit demonstrations, we provide a few hundred demonstrations, recognizing the increased complexity of MoMa.
+MoMa encompasses not only higher-dimensional observation, state, and action spaces but also significantly larger and more varied task spaces.
 This deliberate dataset size ensures sufficient diversity for learning while maintaining a manageable scale that avoids the excessive computational and collection costs of unrealistically large datasets.
@@
 The agent must account for spatial relationships, sequential task dependencies, and the compounding effects of errors over time.
-Error propagation is a critical challenge across many learning paradigms, exemplified by covariate shift in imitation learning.
+Error propagation is a critical challenge across many learning paradigms, exemplified by covariate shift in BC.
 For instance, navigating to a distant room to retrieve an object and returning it to another location involves traversing obstacles and staircases while adhering to task objectives and maintaining situational awareness.
@@
 Additionally, we benchmark a neuro-symbolic system that integrates foundation models with task and motion planning (TAMP).
-This system, evaluated zero-shot without requiring training, investigates whether non-IL methods can achieve greater data efficiency while achieving comparable or superior performance, offering guidance for future development of data-efficient MoMa methodologies.
+This system, evaluated zero-shot without requiring training, investigates whether non-BC methods can achieve greater data efficiency while achieving comparable or superior performance, offering guidance for future development of data-efficient MoMa methodologies.
+Finally, we assess the performance of a random agent, wherein actions are uniformly sampled to establish a lower-bound baseline.
 RT-1.
@@
 To adapt RT-1 for the ManipulaTHOR robot, we modified its action space to align with the robot’s capabilities.
-For task navigation, the robot base operates using discrete actions, while continuous deltas are utilized for body rotation and manipulation to achieve precise control.
-Predicting deltas between timesteps rather than absolute states improved learning stability and control smoothness, while the discretization of continuous action outputs into 256 uniformly distributed bins enabled effective representation and improved learning.
-MotionGlot.
-MotionGlot (MG) [25] is designed to address the problem of varying action and observation dimensions in multi-embodiment models [5, 26].
-Leveraging training principles from language models (LLMs), MG proposes a training procedure for motion generation across multiple embodiments as a next-token prediction problem.
-Here we adapt MG for the MoMa tasks, given a set of demonstration trajectories with image observation ($o_{i}$) action ($a_{i}$) pairs in the form of $\tau=\{(o_{0},a_{0}),(o_{1},a_{1}),...,(o_{T},a_{T}))\}$, MG learns a model to map a given observation to action.
+For task navigation, the robot base operates using discrete actions, while continuous deltas (discretized into 256 bins) are utilized for body rotation and manipulation to achieve precise control.
+MotionGlot-MoMa MG [24] is designed to address the problem of varying action and observation dimensions in multi-embodiment models [25].
+Leveraging training principles from language models, MG proposes a training procedure for motion generation across multiple embodiments as a next-token prediction problem.
+We adapt MotionGlot (MG) for MoMa tasks by leveraging $\lambda$’s demonstrations consisting of paired RGB observations and actions, calling it MG-MoMa.
+MG-MoMa is trained to learn a language-conditioned model that maps those observations to corresponding actions.
 This process involves two stages: First, a VQ-VAE is employed to learn a discrete latent codebook for the images, forming the observation vocabulary, while discrete binning is applied to tokenize the actions.
@@
 This unified vocabulary allows for action generation similar to text generation.
-Finally, the instruction-tuning template from MG is used for text command-based action generation.
+Finally, the instruction-tuning template from MG-MoMa is used for text command-based action generation.
 LIMP.
@@
 Unlike learning-based methods, LIMP requires no training, allowing it to be evaluated zero-shot.
-This enables a comparison of its data efficiency against BC models, providing insight into whether such systems are a promising direction for developing more data-efficient approaches to mobile manipulation.
+This enables a comparison of its data efficiency against BC models, providing insight into whether such systems are a promising direction for developing more data-efficient approaches to MoMa.
 Scene 1 Scene 2 Scene 3 Scene 4 Scene 5 Task Gen
@@
 TABLE II: Data Splits.
-Splitting of the simulation dataset of 521 demonstrations into training, validation, and held-out test subsets for the Scene Generalization and Task Generalization experiments.
+Splitting of the simulation dataset of 521 demonstrations into subsets for the Scene and Task Generalization experiments.
 The “Scenes” row indicates the number of scenes contributing to each subset of demonstrations.
@@
 During each fold, we train and hyperparameter tune a separate model specific to the held-out test scene.
-This process ensures that we assess the model’s ability to generalize to entirely unseen environments while maintaining robust experimental controls.
+This process ensures that we assess the model’s ability to generalize to entirely unseen environments while maintaining experimental controls.
 The rotation of held-out scenes allows for a comprehensive evaluation of scene generalization across the entire dataset.
@@
 Training.
-To benchmark the data efficiency of modern BC methods, we train the RT-1 model from scratch without any pretrained parameters on the simulation tasks.
-This approach evaluates how well these models and their architectures learn from our modestly sized yet realistic $\lambda$ dataset, providing a measure of their data efficiency specifically for diverse and long-horizon MoMa tasks.
-Both models are trained using the scene generalization and task generalization experiments, ensuring that their capabilities are evaluated under varying conditions.
+To benchmark the data efficiency of modern BC methods, we train the RT-1 and MotionGlot models from scratch without any pretrained parameters on the simulation tasks.
+This approach evaluates how well these models and their architectures learn from our modestly sized yet realistic $\lambda$ dataset, providing a measure of their data efficiency specifically for long-horizon MoMa tasks.
+Both models are trained using the scene generalization and task generalization experiments, ensuring their capabilities are evaluated under varying conditions.
 Fine-tuning.
-To further evaluate data efficiency, we fine-tune RT-1 by loading pretrained parameters and testing its performance on the $\lambda$ benchmark simulation tasks.
-We investigate whether pretrained knowledge helps or hurts model generalization on our tasks.
+To further evaluate data efficiency, we fine-tune RT-1 by loading pretrained parameters and testing its performance on the simulation tasks.
 Empirical analysis showed that freezing some model layers led to better performance than not freezing any layers, so this approach was used for the final experiment.
@@
 Scene Generalization.
-The success rates for the Scene Generalization experiment, detailed in Table III, are exceptionally low, with an overall average of just $2.7\%$ across all scenes.
+The success rates for the Scene Generalization experiment, detailed in Table III, are exceptionally low, with an overall average of just $2.7\%$ and $2.4\%$ across all scenes for RT-1 and MG-MoMa respectively.
 None of the individual scenes achieved success rates exceeding $5\%$, indicating significant challenges in generalizing to novel scenes with unseen room layouts, furniture configurations, and object placements for long-horizon MoMa tasks.
-Scenes 1 and 3 exhibited relatively higher success rates compared to the others.
+Both models outperform the random agent, demonstrating their ability to learn from the data and satisfy the minimum performance criterion.
+Scenes 1 and 3 exhibited relatively higher success rates compared to the others for RT-1.
 This can be attributed to a relatively high proportion of evaluation tasks in these scenes involving objects that were positioned directly in front of or adjacent to the robot’s initial spawn location, reducing the complexity of navigation and manipulation required.
-In contrast, Scene 1 exhibited the lowest performance, likely due to the robot spawning in an enclosed and isolated room where many tasks necessitated navigating out of the room to retrieve the target object, significantly increasing task difficulty.
+In contrast, Scene 1 for RT-1 exhibited the lowest performance, likely due to the robot spawning in an enclosed and isolated room where many tasks necessitated navigating out of the room to retrieve the target object, significantly increasing task difficulty.
+Interestingly, for MG-MoMa, Scenes 1 and 3 were the exact opposite of RT-1, with those being the lowest performers.
+It is unclear why that is the case, but it could be that MG-MoMa fails to recognize nearby objects.
+Similarly, Mg-MoMa’s best performers are RT-1’s worst, Scenes 2 and 5.
+Overall, MG-MoMa demonstrates performance that is near RT-1, which is noteworthy given that Motionglot was not originally intended for mobile manipulation tasks.
 Task Generalization.
-The success rates for the Task Generalization experiment for RT-1 and MotionGlot, as presented in Table IV, are low overall, though they are slightly higher compared to the average performance in the Scene Generalization experiment.
+The success rates, presented in Table IV, for the Task Generalization experiment for RT-1 and Mg-MoMa, are low overall, though they are slightly higher compared to the average performance in the Scene Generalization experiment.
 This improvement is likely attributed to the model having access to all the scenes during training, enabling it to better generalize to new tasks within already familiar environments.
-Additionally, while it is unsurprising that RT-1 outperformed MotionGlot, it is notable that the margin was relatively small, particularly given that MotionGlot was originally designed for motion generation rather than specifically for mobile manipulation.
+Additionally, while it is unsurprising that RT-1 outperformed MG-MoMa on unseen demonstrations, it is notable that the two models performed comparably on seen demonstrations, and the margin of RT-1’s outperformance was relatively small.
+This is remarkable, particularly given that MotionGlot was originally designed for cross-embodied motion generation rather than specifically for mobile manipulation.
 We also evaluated on seen tasks, which were uniformly sampled from the training data to match the size of the unseen data from the test split to measure performance on previously encountered tasks.
@@
 Task Generalization with Fine-Tuning.
-To evaluate whether pretrained parameters improve generalization on our benchmark tasks, we fine-tuned RT-1 with some of its layers frozen and tested it on both seen and unseen tasks.
 Interestingly, the unseen task performance was marginally lower compared to the fully trained model’s performance on unseen tasks, as seen in Table IV.
-This result suggests that the pretrained weights may have skewed the model’s representation toward the original training tasks and away from effectively adapting to the unique challenges of our benchmark, especially since the original was trained in the real world
+This result suggests the pretrained weights may have skewed the model’s representation toward the original training tasks and away from effectively adapting to the unique challenges of our benchmark, especially since the original was trained in the real world.
 LIMP.
-The results presented thus far indicate that learning-based models, such as RT-1 and MotionGlot, face significant challenges in addressing the complexities of long-horizon MoMa tasks.
+The results presented thus far indicate that learning-based models, face significant challenges in addressing the complexities of long-horizon MoMa tasks.
 To explore an alternative paradigm, we benchmarked the LIMP system on the Task Generalization experiment.
@@
 For these simulation-designed tasks, real-world setups were constructed to approximate their simulated counterparts.
-This setup enabled us to evaluate LIMP’s versatility across both task types, providing insights into its adaptability to different environmental contexts.
-As shown in Table IV, LIMP achieved a relatively impressive average success rate of $44.4\%$ on all novel tasks, vastly outperforming RT-1 without requiring any training, making it relatively more data efficient.
+This enabled us to evaluate LIMP’s versatility across both task types, providing insights into its adaptability to different environmental contexts.
+As shown in Table IV, LIMP achieved a relatively impressive average success rate of $44.4\%$ on unseen tasks, vastly outperforming RT-1 and MG-MoMa without any training, making it relatively more data efficient.
 One key factor contributing to LIMP’s success is its use of pre-generated maps of the environments.
-These maps provide the robot with precise information about the absolute and relative locations of objects and regions, eliminating the need to search and enabling it to interpret the fine-grained semantics specified in language-conditioned tasks.
+They provide the robot with precise information about the absolute and relative locations of objects and regions, eliminating the need to search and enabling it to interpret the fine-grained semantics specified in language-conditioned tasks.
 Additionally, the maps allow the robot to utilize robust motion planning algorithms for navigation, which are less susceptible to covariate shift compared to the mapless autoregressive action decoding used in BC models like RT-1.
-This robustness is particularly critical for the long-horizon navigation aspect of the tasks, where the inherent large observation and state spaces exacerbate the challenges faced by learning-based methods.
+This robustness is critical for the long-horizon navigation aspect of the tasks, where the inherent large observation and state spaces exacerbate the challenges faced by learning-based methods.
 Another significant advantage of LIMP is its incorporation of multimodal foundation models, which endow it with extensive prior knowledge to reason about task requirements and perceptual inputs in a zero-shot manner.
@@
 This evaluates whether these models can effectively utilize limited data or require substantially more, directly testing their data efficiency and scalability for long-horizon MoMa tasks.
-The results from the Task Generalization setup, summarized in Table V, surprisingly reveal no clear upward trend in performance gains despite progressively increasing the training data for RT-1 and hyperparameter tuning per subset.
+The ablation results from the Task Generalization setup, summarized in Table V, surprisingly reveal no clear upward trend in performance gains despite progressively increasing the training data for RT-1 and hyperparameter tuning per subset.
 We attribute part of the variation to the inherent complexity of these long-horizon MoMa tasks and the variability in the sampled subsets, where smaller subsets can incidentally include a higher proportion of comparatively easier scenarios.
@@
 RT-1 $4.9\pm 9.9$ $0.78\pm 4.4$ $4.3\pm 9.4$ $2.0\pm 6.7$ $1.4\pm 5.8$ $2.7$
+MG-MoMa $1.9\pm 5.7$ $3.7\pm 8.9$ $0.8\pm 5.7$ $2.0\pm 6.7$ $3.8\pm 9.0$ $2.44$
+Random $0.7\pm 4.1$ $0.5\pm 3.6$ $0.3\pm 2.5$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.3$
 TABLE III: Scene Generalization results.
@@
 RT-1 $4.6\pm 9.7$ $5.1\pm 10.1$ $4.6\pm 9.7$ $4.6\pm 9.7$
-MotionGlot $3.24\pm 8.4$ $2.4\pm 7.2$
+MG-MoMa $4.6\pm 9.7$ $4.2\pm 9.3$
 LIMP $44.4\pm 49.7^{*}$
+Random $0.5\pm 3.4$ $0.5\pm 3.4$
 TABLE IV: Task Generalization results.
@@
 Furthermore, since RT-1’s $100\%$ performance is somewhat better than the LSTM variant’s, and its RoC is slightly higher, it indicates marginally better data efficiency, further supporting the argument that architecture is unlikely the primary cause of RT-1’s data inefficiency observed in the main experiments.
-Since architecture does not appear to be the primary issue, it is plausible that these long-horizon indoor MoMa tasks are inherently complex and challenging for learning-based methods.
-Identifying the exact causes of data inefficiency in these models requires further investigation, which is beyond the scope of this paper and is left for the community to explore in future work.
+Since architecture does not appear to be the primary issue, it is plausible that these long-horizon indoor MoMa tasks are simply inherently too complex and challenging for learning-based methods.
+Identifying the exact causes of data inefficiency in these models requires further investigation, which is beyond the scope of this work and is left for the community to explore in future work.
 VI Conclusion
-We introduce LAMBDA ($\lambda$), a challenging benchmark consisting of a realistic amount, 571, of language-conditioned, long-horizon mobile manipulation tasks, spanning room-to-room and floor-to-floor navigation in indoor environments.
+We introduce LAMBDA ($\lambda$), a challenging benchmark consisting of a realistic amount, 571, of language-conditioned, long-horizon MoMa tasks, spanning room-to-room and floor-to-floor navigation in indoor environments.
 Each task is paired with human-collected expert demonstrations from both simulated and real-world settings.
-To evaluate the benchmark, we tested multiple models: RT-1, MotionGlot, and LIMP, to assess their performance and data efficiency on these complex tasks.
-Learning-based models (RT-1 and MotionGlot) demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to $\lambda$’s long-horizon tasks.
-In contrast, the neuro-symbolic system LIMP significantly outperformed RT-1 and MotionGlot, showcasing its potential as a promising direction for the development of future mobile manipulation systems capable of addressing the challenges posed by $\lambda$’s tasks.
+To evaluate the benchmark, we tested multiple models to assess their performance and data efficiency on these complex tasks.
+Learning-based models, RT-1 and MotionGlot, demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to $\lambda$’s long-horizon tasks.
+In contrast, the neuro-symbolic system LIMP significantly outperformed the BC models, showcasing its potential as a promising direction for the development of future MoMa systems capable of addressing the challenges posed by $\lambda$’s tasks.
 We hope that this benchmark will serve as a valuable resource for the robotics community, fostering the development of more advanced models and systems that can efficiently address daily household and workplace tasks.
@@
 Future work could also benchmark reinforcement learning models, which, while not the focus of this study, are known for their sample inefficiency.
-Evaluating their performance on $\lambda$’s sparse rewards would offer additional insights into their suitability for long-horizon mobile manipulation tasks and broaden the scope of the benchmark.
+Evaluating their performance on $\lambda$’s sparse rewards would offer additional insights into their suitability for long-horizon MoMa tasks and broaden the scope of the benchmark.
 Model $25\%$ $50\%$ $75\%$ $100\%$ Avg RoC
@@
 LSTM $8.3\pm 11.8$ $2.7\pm 7.7$ $5.4\pm 10.3$ $1.4\pm 5.7$ $-0.092$
-TABLE V: Ablations results.
-Success rates for Task Generalization on growing subsets of the simulation dataset for both RT-1 and its LSTM variant.
-The computation is the same as Table III.
-“Avg RoC” indicates the average rate of change between subset results.
+TABLE V: Ablations results. Success rates for Task Generalization on growing subsets of the simulation dataset. “Avg RoC” indicates the average rate of change between subset results.